ydy=4x(y^2+1)^1/2 dx y(0)=1


1
Expert's answer
2021-03-05T07:26:39-0500

"ydy = 4x{\left( {{y^2} + 1} \right)^{\frac{1}{2}}}dx,\,\,\,y(0) = 1"


"\frac{{ydy}}{{{{\left( {{y^2} + 1} \right)}^{\frac{1}{2}}}}} = 4xdx"


"\frac{1}{2}\frac{{d\left( {{y^2} + 1} \right)}}{{{{\left( {{y^2} + 1} \right)}^{\frac{1}{2}}}}} = 4xdx"


"\sqrt {{y^2} + 1} = 2{x^2} + C"


"\sqrt {{y^2} + 1} - 2{x^2} = C"


"y(0) = 1 \Rightarrow \sqrt {{1^2} + 1} - 0 = C \Rightarrow C = \sqrt 2"


Answer: "\sqrt {{y^2} + 1} - 2{x^2} = \sqrt 2"